OpenStudy (anonymous):

PLZZ HELP! Are 3x+33 and 3 equivalent? ----- x+11 Using complete sentences explain why or why not.

OpenStudy (anonymous):

Yess

OpenStudy (lalaly):

factor out the 3 from the nominator

OpenStudy (anonymous):

\[\frac{ 3x+33 }{ x+11 }=\frac{ 3(x+11) }{ x+11 }=3\] cancel out the x+11 on top and bottom and you remain with 3